OpenStudy (anonymous):

Is the following problem simplified correctly? (5x+7)(5x-7)=25x2+70x-49

OpenStudy (anonymous):

no. (a+b)(a-b) = a^2 - b^2. there is no ab term.

OpenStudy (anonymous):

@bdboy, same as before... when you have the product of a sum and difference of binomials you have a factored form of a difference of two squares, that is, the product is a difference of two squares (se the right side of dhatraditya's formula?)

OpenStudy (anonymous):

in your problem a=5x and b=7

OpenStudy (anonymous):

yes, (a+b)(a-b) = a^2 +ba -ab - b^2 = a^2 +ab-ab-b^2 = a^2-b^2

OpenStudy (anonymous):

when you have the product of a sum and difference of binomials, the middle two terms of the FOIL will always cancel each other out

OpenStudy (anonymous):

thanks a ton i really appreciate you not just giving me the answers but actually explaing how to get them which helps me understand it so much more...

OpenStudy (anonymous):

you need to multiply 25 and 2 then subtract 49

OpenStudy (anonymous):

ok, what's the product

OpenStudy (anonymous):

do you want me to show you

OpenStudy (anonymous):

it think i totally just screwed it up.. somehow i got 1 but thats not correct

OpenStudy (anonymous):

please that would be most helpful

OpenStudy (anonymous):

that is the correct answer for zac's question

OpenStudy (anonymous):

once you realize you have a difference of squares you just write\[(5x-7)(5x+7)=(5x)^2-7^2=25x^2-49\]with practice you would most likely do the middle step as a mental calculation

OpenStudy (anonymous):

notice on the foil how the middle two terms cancel out\[(5x-7)(5x+7)=25x^2+35x-35x+49=25x^2-49\]

OpenStudy (anonymous):

if I may, with mandolino's permission, \[(a+b)(a-b) = a(a-b) + b(a-b) = a^2 -ab + ba -b^2\] \[= a^2 - ab + ab -b^2 = a^2 - b^2\]

OpenStudy (anonymous):

note that i just noticed a sign error in my previous post: the +49 in the middle step should be -49; the final answer is correct

OpenStudy (anonymous):

@dhatraditya, i've just been playing off your posts

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

bdboy try this one: (2x-3)(2x+3)

OpenStudy (anonymous):

i'm going to give the answer in my next post, so don't look if you are trying it on your own!

OpenStudy (anonymous):

\[(a-b)(a+b)=a^2-b^2\]in my example a=2x and b=3\[(2x-3)(2x+3)=(2x)^2-3^2=4x^2-9\]

OpenStudy (anonymous):

can you try one more except let me put the answer down